Accelerator Update for September 24, 2008 - September 26, 2008
Wednesday Sept 24
The day shift began with Operations monitoring the Tevatron (TeV) store 6442 (79.96E30) and the Recycler stash (251.7E10), with the Antiproton Source (Pbar) stacking (24.1E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE, NuMI, and the Meson MTest experiment taking beam, and with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams.
Operations reported at 11:05 AM that they had restarted a Booster vacuum pump (IP132), a MI vacuum pump (IP307A), and a MI-8 transfer line vacuum pump (IP827).
The Neutron Therapy Facility reported at 11:07 AM that they had treated their last patient for the day.
Operations terminated store 6442 at 2:20 PM.
Operations put the TeV into shot setup at 2:31 PM.
Operations reported at 4:01 PM that the Recycler had stashed antiprotons five times during the day shift.
A MI expert began a 30-minute study period at 4:53 PM.
Operations disabled the MTest beam at 5:58 PM.
Operations put the TeV through a dry squeeze at 7:02 PM.
At 8:21 PM, Operations reported receiving many spurious FIRUS readings that were not believable. Operations contacted a FIRUS expert.
At 8:47 PM, the TeV aborted due to bad beam positions at sector F4.
At 8:49 PM, the MI RF south anode power supply tripped off. Everything reset.
Cryo system experts reported at 9:06 PM that they had completed the TeV sector A3 flywheel change out.
An RF expert disabled MIRF1 at 10:20 PM due to a modulator problem.
At 10:25 PM, LRF4 tripped off. It reset.
At 10:47 PM, BRF8 tripped off. It reset.
Operations put the TeV through a wet squeeze around 11:12 PM.
At 11:44 PM, MIRF15 tripped off. Operations reduced the NuMI intensity and reset the station.
Operations checked the TeV coalescing at 11:51 PM and reported that it would get by.
Operations reported at midnight that the Recycler had stashed five times during the evening shift.
Thursday Sept 25
The midnight shift began with the TeV in a wet squeeze, with Operations monitoring the stash (200.9E10), with the Pbar stacking (26.2E10), and with the MiniBooNE and NuMI experiments taking beam.
Operations put the TeV into a shot setup at 12:48 AM.
Operations established store 6447 at 2:42 AM with an initial luminosity of 196.8E30.
Operations established beam to MTest at 4:38 AM.
At 5:39 AM, LRF4 tripped off for the first of two times in 20 minutes. It reset.
Operations reported at 7:55 AM that the Recycler had stashed four times during the midnight shift.
A Pbar expert reported at 8:13 AM that he was adjusting the target position.
At 8:20 AM, a Linac Klystron RF station (KRF2) tripped off. It reset.
At 10:06 AM, TRF6 tripped off. It reset.
Operations prepared the Pbar AP0 vault for access at 10:26 AM.
Operations put Pbar into standby at 10:32 AM.
At 10:33 AM, the MI Power Supply interlock Loop (MIPSLoop) tripped off due to a volts-to-ground indications on the MI-10 and MI-20 lower busses. Operations contacted an expert who managed to reset the trips. He then began investigating the cause.
Operations reported at 10:33 AM that all the MIRF stations were off so that stations MIRF1 and MIRF15 could be repaired.
AP0 came out of access at 10:57 AM.
MI experts reported at 11:05 AM that all RF work had been completed.
Pbar resumed stacking at 11:07 AM.
At 12:53 PM, KRF4 tripped off for the first of four times in 13 minutes. Operations contacted a Linac expert.
Operators completed searching and securing MTA-MuCool at 2:03 PM.
At 2:47 PM, LRF5 tripped off. It reset.
Operations terminated store 6447 at 3:07 PM.
Operations reported that TeV pulsed power supply thermal scans began at 3:09 PM. This is an ongoing check. (Later, they reported that the knife switch at A3 had a high temperature.)
At 3:33 PM, LRF2 tripped off. Experts had the station back on about 25 minutes later.
Operations reported at 4 PM that the Recycler had stashed four times during the day shift.
Operations disabled the MTest beam at 4:18 PM.
A TeV RF expert reported 4:22 PM that TRF6 had failed and needed to be replaced. He said the replacement would take about 30 minutes.
Operations put the TeV into shot setup at 4:28 PM.
Operations reported at 5:41 PM that a Booster expert was tuning the accelerator.
At 5:54 PM, LRF4 tripped off. It reset.
Operations established store 6448 at 6:01 PM with an initial luminosity of 268.9E30.
CDF reported at 9:01 PM that the Tevatron Electron Lens (TEL) had tripped off. Operators reset it.
At 9:58 PM, LRF2 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Friday Sept 26
The midnight shift began with Operations monitoring store 6448 (121.26E30) and the stash (174.2E10), with Pbar stacking (20.8E10), and with the MiniBooNE and NuMI experiments taking beam.
Operations reported at 12:14 AM that there was a small drop in luminosity early on the evening shift. Both CDF and D0 saw the drop, but nothing alarmed.
Operations established beam to MTest at 4:13 AM.
Operations reported at 7:56 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 09/29/2008 email Fermilab|
|Security, Privacy, Legal|